Calcium carbonate reacts with aqueous HCl to give CaCl2 and CO2 according to the reaction, CaCO3 (s) + 2 HCl (aq) ? CaCl2 (aq) + CO2(g) + H2O(l) What mass of CaCO3 is required to react completely with 25 mL of 0.75 M HCl?

Grade:12

4 Answers

Gaurav
9 years ago
0.75 M solution of HCl = 0.75 moles of HCl in 1 L water
Number of moles of HCl in 25mL of 0.75 M solution = (0.75/1000)25 =0.019 moles
From equation it is clear that one mole CaCO3 reacts with two moles of HCl
So,
Number of moles of CaCO3 required to react with 0.019 moles of HCl = (1/2)0.019 = 0.001
Mass of 0.001 mole CaCO3 = (0.001 )(Molar mass of CaCO3) =(0.001 )( 100) = 0.1g
Mihir
6 years ago
CaCO3-1 mol = 100gHCl-2 mol = 73g0.75M HCl - 0.75 moles of HCl in 1000mlIn 25ml = 0.75 ×25/1000 = 0.01875 molSince 1 mol HCl reacts with 2 mol CaCO3 0.01875molHCl = 0.01875/2 = 0.009375 mol CaCO3In grams= 0.009375 × molar mass of CaCO3= 0.9375g
Tanu
5 years ago
No. Of moles of HCL used = Mass× Volume÷1000=  0.75×25÷1000
                         =0.01875
                         =0.019
No.of moles of CaCO3 to react with HCL=1÷2 ×no.of moles of HCL   
=1÷2×0.019÷1000
=9.5×10 to the power -3
Mass of CaCO3 = no. Of moles of CaCO3 × molar mass
                            = 9.5×10to the power -3 × 100
                            = 0.95 g
 
 
 
 
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

0.75 M solution of HCl = 0.75 moles of HCl in 1 L water
Number of moles of HCl in 25mL of 0.75 M solution
= (0.75/1000)25
=0.019 moles
From equation it is clear that one mole CaCO3 reacts with two moles of HCl
So,
Number of moles of CaCO3 required to react with 0.019 moles of HCl
= (1/2)0.019
= 0.001 Mass of 0.001 mole CaCO3
= (0.001 )(Molar mass of CaCO3)
=(0.001 )( 100)
= 0.1g

Thanks and Regards